Ethylene Glycol is used in automobile antifreeze and preparation of polyester fiber. Ethylene glycol is produced in the following reaction:


C2H2O  + H2O ----> C2H6O2  


If 90 g of water are used completely in the reaction, how many moles of ethylene glycol are produced? How many grams is this equivalent to?



1
Expert's answer
2021-01-07T05:32:52-0500

Assuming that your equation is wrong (that you meant C2H4O and not C2H2O as the reagent).


Because Ethylene glycol is produced from ethene via an intermediate ethylene oxide (C2H4O). It is ethylene oxide that is reacted with water to form ethylene glycol as shown below;


C2H4O + H2O "\implies" C2H6O2


Mole ratio of all species involved is 1:1:1


Moles of water in 90g

"Moles = \dfrac{mass}{MM} = \dfrac{90g}{18g/mol} = 5mol"



Moles of ethylene glycol according to the equation


Moles of C2H6O2 ="\dfrac{1}{1}x 5mol = 5 mol"


Mass of C2H6O2 = Moles x MM of C2H6O2

= 5 mol x 62 g/mol

= 310g